OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=6/7,(4,-6) The equation of the line is y=_________ Can someone please tell me what Y equals?

OpenStudy (caozeyuan):

y--6=6/7*(x-4) y+6=6/7x-24/7

OpenStudy (anonymous):

This still doesn't seem complete. It asks what Y equals. Am I missing something?

OpenStudy (shamim):

ok

OpenStudy (shamim):

i m trying

OpenStudy (anonymous):

Thank you

OpenStudy (caozeyuan):

if you just pu t +6 to the RHS, you'll find the and

OpenStudy (shamim):

the formula is\[y-y _{1}=m(x-x _{1})\]

OpenStudy (shamim):

here \[x _{1}=4,y _{1}=-6\]

OpenStudy (caozeyuan):

y=(6/7)*x-66/7

OpenStudy (shamim):

so\[y-(-6)=\frac{ 6 }{ 7 }(x-4)\]

OpenStudy (shamim):

\[y+6=\frac{ 6 }{ 7 }(x-4)\]

OpenStudy (shamim):

\[y=\frac{ 6 }{ 7 }x-\frac{ 6 }{ 7 } \times 4-6\]

OpenStudy (shamim):

is it clear

OpenStudy (anonymous):

not at all! I'm sorry!

OpenStudy (shamim):

why

OpenStudy (shamim):

if i say the equation is\[6x-7y=66\]

OpenStudy (anonymous):

On the test I just took with the same question, different numbers: m=3/4,(9,-4) The correct answer, which is what I missed, was y=3/4x-43/4

OpenStudy (anonymous):

I just can't figure out this problem. What does y equal?

OpenStudy (caozeyuan):